OpenStudy (anonymous):

Solve the equation: 7x = 42

OpenStudy (samsan9):

6....

OpenStudy (anonymous):

just divide both sides by 7 x=42/7=6 =>x=6

OpenStudy (anonymous):

yup just divide 42 by 7

OpenStudy (solomonzelman):

how about 7x² = -42 ? :)

OpenStudy (anonymous):

I typed the wrong question on this one, sorry guys. I knew that answer.

OpenStudy (anonymous):

7x = 42 (divide both sides by 7) 7x/7 = 42/7 x = 6

OpenStudy (anonymous):

So the answer is x = 6